spherical metal shell A of radius Ra and solid metal sphere B of radius Rb are given charge +Q and are then connected by a thin metallic wire.
Which of the following are true?
A.Qa>Qb
B.EA
c.SIGMA A/SIGMA B =Rb/Ra
sigma is surface charge density
Ra>Rb
 
 

Grade:12

1 Answers

Khimraj
6 years ago
Both sphere are connected by thin metallic wire. So there surface will be at same potential . then
KQa/Ra =KQb/Rb
then Qa/Ra =Qb/Rb
(Qa/4πRa2)4πRa = (Qb/4πRb2)4πRb
Then  (\sigmaa)Ra = (\sigmab)Rb
So \sigmaa/\sigmab = Rb/Ra.
Hope it clears. If you like answer then please approve it.